# A GLOBAL MINIMUM TAX FOR LARGE FIRMS ONLY:

### v6

@@ Appendix E Welfare Effects of Increasing the GMT Coverage Rate
Proposition E.1
(Increasing in the GMT coverage rate $\phi$)
@@
In summary, the welfare effects of an increase in the GMT coverage rate $\phi$ are similar to those of an increase in the GMT rate $t_{M}$ (Proposition 3). A higher $\phi$ has opposing welfare effects in non-haven and haven countries in Regimes 2 to 4, but not in Regime 1. One notable difference from the effects of a higher $t_{M}$ is that a higher $\phi$ can never harm the non-haven country and the world, even if it induces a regime switch.
@@
This completes the proof of Proposition E.1. ∎
Appendix F Supplementary Figures
Figure F.1 shows the welfare levels of the non-haven (panel (a)) and the sum of the havens (panel (b)), and the shifted profits for different GMT rates $t_{M}$.
The tax-competition equilibrium outcomes under partial GMT coverage (solid curves) are compared with those under full GMT coverage (dashed curves). The equilibrium tax rates under full GMT are as follows.
Let $G_{n}^{f}$ and $G_{h}^{f}$ be the equilibrium payoffs of the non-haven and the haven under full GMT. In Regime 0 ($t_{M}\in[0,t_{M}^{0})$), the equilibrium tax rates are identical under full and partial GMT coverage, as neither country is constrained by the GMT.
Fig. F.1. Equilibrium Welfare and Shifted Profits.
Notes: The figure shows the equilibrium levels of welfare in the non-haven ($G_{n}$) in panel (a) and the sum of the $H=40$ havens ($\sum_{h=1}^{H}G_{h}$) in (b), and the aggregate shifted profits ($\sum_{h=1}^{H}\theta_{h}\Pi$) in (c), for different GMT rates $t_{M}$. The solid curves in each panel are the equilibrium values under the partial GMT ($\phi\in(3/4,1)$); the dashed curves are those under the full GMT ($\phi=1$); horizontal lines are those under no GMT (Regime 0). While the levels of welfare, shifted profits and regime-switching points vary with parameter values, the qualitative patterns remain unchanged.

### v7

@@ D.2 Welfare effects of increasing the GMT coverage rate
Proposition D.2. (Increasing in the GMT coverage rate $\phi$)
@@
In summary, the welfare effects of an increase in the GMT coverage rate $\phi$ are similar to those of an increase in the GMT rate $t_{M}$ (Proposition C.1). A higher $\phi$ has opposing welfare effects in non-haven and haven countries in Regimes 2 to 4, but not in Regime 1. One notable difference from the effects of a higher $t_{M}$ is that a higher $\phi$ can never harm the non-haven country and the world, even if it induces a regime switch.
@@
This completes the proof of Proposition C.2. ∎
Appendix E Extensions of the basic model
E.1 Decentralised decision-making by tax havens
In the text, we assumed that a group of tax-haven countries collectively decide whether to adopt a single GMT rate or to split their tax rate. Here, we consider an alternative, decentralised decision-making process in the first stage of our game, where each tax haven individually decides whether to split, given the decisions of the others. In the following, we focus on $t_{M}\in[t_{M}^{0},t_{M}^{2}]$ and impose two assumptions. First, countries can split their tax rates only if they are subject to the GMT. This assumption reflects the fact that preferential tax regimes for specific tax bases are not permitted under the OECD’s action plan to address base erosion and profit shifting (OECD, 2019). Second, each haven forms the same belief about how the other havens and the non-haven choose their tax rates. This allows us to focus on the problem of a single haven.
$\begin{split}&t_{n}(H_{M}-1)=\frac{2(\lambda-1)\left[t_{M}\left\{H\phi+(H_{M}-1)(1-\phi)\right\}+\delta\right]}{(\lambda-1)[\phi(H-H_{M}+1)+H+H_{M}-1]+2H\lambda},\\ &t_{h}^{2}(H_{M}-1)=\begin{cases}t_{n}(H_{M}-1)/2&\text{for}\ \pi\in[\underline{\pi},\pi_{M})\\ t_{M}&\text{for}\ \pi\in[\pi_{M},\infty)\end{cases},\end{split}$ (A.7)
The difference between the haven $h$’s payoffs in the two cases is
$\Delta G_{h}^{12}(H_{M})\equiv G_{h}^{1}(H_{M})-G_{h}^{2}(H_{M}-1)$
where
$\Theta(H_{M})\equiv AH_{M}+B,$
where
$E\equiv\delta H(2\lambda-1)(\lambda-1)>0,$
We can check $t_{M}^{0}<t_{M}^{a}<t_{M}^{b}<t_{M}^{1}$:
$t_{M}^{b}-t_{M}^{a}=\frac{\delta(H-1)(1-\phi)^{2}(\lambda-1)^{4}}{DF}>0,$
which is indeed positive.
The discussion is summarised as follows. If $t_{M}\in[t_{M}^{0},t_{M}^{a}]$, the non-haven chooses a single non-GMT rate, and each haven chooses a single GMT rate, as in Regime 1 in the text (Proposition 1(i)). If $t_{M}\in(t_{M}^{b},t_{M}^{2}]$, the non-haven sets a single non-GMT rate, and each haven splits its tax rate and chooses a GMT rate for large MNEs, but a lower rate than the GMT for small MNEs, as in Regime 2 in the text (Proposition 1(ii)). If $t_{M}\in[t_{M}^{a},t_{M}^{b})$, both Regimes 1 and 2 can be equilibria and either may emerge. If $t_{M}\in[t_{M}^{a},t_{M}^{b})$ and in addition $H_{M}$ coincides with $H_{M}^{*}$ at which $\Delta G_{h}^{12}(H_{M}=H_{M}^{*})=0$, the non-haven sets a single non-GMT rate ($t_{n}(H_{M}^{*})$ in (A.7)), and there are $H_{M}^{*}$ havens setting a single GMT rate $t_{M}$ and $H-H_{M}^{*}$ havens splitting their tax rates ($t_{h}^{2}(H_{M}^{*})$ in (A.7)).
A clear difference from Proposition 1 is that the range where only Regime 1 occurs is smaller ($t_{M}^{0},t_{M}^{a})\subset[t_{M}^{0},t_{M}^{1})$) when havens make their first stage decision in a decentralised manner, implying that they are less likely to commit to a single GMT rate. This is intuitive because the commitment of a haven induces the non-haven to raise its tax rate less than the simultaneous commitment of many havens.
E.2 Real responses of MNEs’ profits to taxes
As described in the text, we allow the pre-tax profits of MNEs, $\pi$, to vary in taxes. Specifically, we assume $\pi=\pi^{b}/(1+t_{n})$, where the baseline pre-tax profit, $\pi^{b}\in[\underline{\pi},\infty)$, follows a cumulative distribution function $F(\pi^{b})$. The total pre-tax profits are
$\Pi=\int_{\underline{\pi}}^{\infty}\frac{\pi^{b}}{1+t_{n}}dF=\frac{\Pi^{b}}{1+t_{n}},$
where $\Pi^{b}\equiv\int_{\underline{\pi}}^{\infty}\pi^{b}dF$ is the total baseline pre-tax profits. We experimented with an alternative specification, $\pi=\pi^{b}(1-t_{n})$, and found that the quantitative implications of introducing a 15% GMT rate are similar to those here.
We adopt the calibration strategy described in Section 5 of the text. Specifically, we calibrate the unconstrained model without GMT (Regime 0) to match key data on international profit shifting. We set the GMT coverage rate to $\phi=0.9$ and the number of haven countries to $H=40$, as in Section 5. We jointly choose the MVPF $\lambda$ and the profit-shifting cost parameter $\delta$ to minimize the sum of squared distances between the model-implied non-haven tax rate and shifted profits and their empirical counterparts, $(t_{n}^{0}-0.186)^{2}+(\sum_{h}\theta_{h}^{0}-0.209)^{2}$, where
$t_{n}^{0}=\frac{2\left[\sqrt{H\left[(16\delta+9H)\lambda^{2}-(38\delta-6H)\lambda+12\delta+H\right]}-H(3\lambda-1)\right]}{H(8\lambda-3)},\qquad t_{h}^{0}=\frac{t_{n}^{0}}{2},$
$\theta_{h}^{0}=(t_{n}^{0}-t_{h}^{0})/\delta.$
The calibration results are given in Table E.1 and the quantitative effects of GMT reforms in Table E.2. The results are similar to those in Section 5. As discussed in the text, the only notable difference is the higher calibrated value of the MVPF, $\lambda=6.0$ (vs. $\lambda=2.1$ in Table 2). Otherwise, calibrated values and the magnitude of the quantitative findings are similar to those in the basic model.
Table E.1
Calibration of the model with MNEs’ pre-tax profits dependent on taxes.
Parameter Source
$\phi=0.9$ Profit share of MNEs covered by GMT OECD (2020), Orbis
$H=40$ Number of tax havens Tørsløv et al. (2023)
$\Pi^{b}=$5,480bUSD Baseline pre-tax profits of MNEs Calibrated
$\lambda=6.0$ Marginal valuation of public funds Calibrated
$\delta=17.7$ Cost of profit shifting Calibrated
Targeted moment Model Data Source
Pre-tax profits of MNEs: $\Pi=\Pi^{b}/(1+t_{n}^{0})$ 4,623bUSD 4,623bUSD Tørsløv et al. (2023), CbCR
Corp. tax rate of non-haven: $t_{n}^{0}$ $\% 18.618.6\%$ Tørsløv et al. (2023)
Share of shifted profits: $\sum_{h}\theta_{h}^{0}$ $\% 20.920.9\%$ Tørsløv et al. (2023), CbCR
Non-targeted moment Model Data Source
Corp. tax rate of haven: $t_{h}^{0}$ $9.3\%$ $13.7\%$ Tørsløv et al. (2023)
Non-haven’s revenue loss: $t_{n}^{0}\sum_{h}\theta_{h}^{0}\Pi$ 180180bUSD 247247bUSD Tørsløv et al. (2023)
Notes: Data are from 2019. The definitions of non-haven and haven countries follow Tørsløv et al. (2023) and the data is from the authors’ website: https://missingprofits.world (corresponding figures in the spreadsheets of ‘Table 1’ and ‘Table U1’ in the excel file ‘1975-2019 updated estimates: Tables‘). The non-haven countries are $30$ OECD countries, $7$ major developing countries and the rest of the world. There are $40$ haven countries across the world. The corporate tax rates for the representative non-haven and haven countries are calculated as the GDP-weighted averages of their respective effective tax rates. Tørsløv et al. (2023) report the pre-tax profits of foreign-owned MNEs only, which do not include those of domestically-owned MNEs (see their Section 3.1.2 on p.7). Using the CbCR Statistics of the EU Tax Observatory (https://www.taxobservatory.eu/repository/the-cbcr-explorer/), we compute the pre-tax profit ratio of domestically-owned MNEs to foreign-owned MNEs. With this ratio ($0.785$) and the pre-tax profits of foreign-owned MNEs ($2,590$ billion USD), we obtain the total pre-tax profits of foreign- and domestically-owned MNEs as $(1+0.785)\times 2,590=4,623$ billion USD.
Table E.2
GMT reforms when MNEs’ pre-tax profits depend on taxes.
(a) Gains from a 15% GMT with $\phi=0.9$ Non-haven Haven World
Tax rate after GMT introduction $20.7\%$ for all $15.0\%$ for all —–
Welfare change in bUSD (% change) $108.9$ ($8.3\%$) 2.5-2.5 ($2.8\%$) $106.3$ ($7.6\%$)
Revenue change in bUSD (% change) $140.7$ ($20.8\%$) 2.5-2.5 ($2.8\%$) $138.2$ ($18.0\%$)
Profit change in bUSD (% change) $386.9$ $(39.9\%)$ 386.9-386.9 $(-39.9\%)$ 0 $(0\%)$
(b) Gains from a 16% GMT with $\phi=0.9$ Non-haven Haven World
Tax rate after GMT introduction $20.8\%$ for all $\begin{cases}$10.4\%$&\text{for small}\\ $16.0\%$&\text{for large}\end{cases}$ —–
Welfare change in bUSD (% change) $118.1$ ($9.0\%$) 7.4-7.4 ($8.3\%$) $110.7$ ($7.9\%$)
Revenue change in bUSD (% change) $152.0$ ($22.4\%$) 7.4-7.4 ($8.3\%$) $144.6$ ($18.8\%$)
Profit change in bUSD (% change) $416.5$ $(43.0\%)$ 416.5-416.5 $(-43.0\%)$ 0 $(0\%)$
[…]
Tax rate after GMT introduction $21.0\%$ for all $16.0\%$ for all —–
Welfare change in bUSD (% change) $129.3$ ($9.9\%$) 7.4-7.4 ($8.3\%$) $121.9$ ($8.7\%$)
Revenue change in bUSD (% change) $166.3$ ($24.5\%$) 7.4-7.4 ($8.3\%$) $158.9$ ($20.7\%$)
Shifting-profits change in bUSD (% change) $453.6$ $(46.8\%)$ 453.6-453.6 $(-46.8\%)$ 0 $(0\%)$
Notes: Panel (a) shows the effects of raising the GMT rate from $t_{M}=0$ to $0.15$ with partial coverage ($\phi=0.9$); panel (b), from $t_{M}=0$ to $0.16$ with $\phi=0.9$; and panel (c), from $t_{M}=0$ to $0.17$ with full coverage ($\phi=1.0$). The tax rates in the initial equilibrium (Regime 0) are $0.186$ in the non-haven and $0.093$ in the haven. The regime-switching rate between Regimes 1 and 2 is 15.9%. Results are based on the calibrated parameter values described in Table E.1. The welfare levels are normalized by the MVPF, $\lambda$. The ‘haven’ here refers to the sum of the $H=40$ haven countries.